OpenStudy (anonymous):

By writing (cos x) in terms of (1/2)x, find an alternative expressions for \[{1-cosx \over 1+ \cos x}.\]

OpenStudy (anonymous):

(tan(x/2))^2

OpenStudy (anonymous):

How?

OpenStudy (campbell_st):

things you need to know \[\cos \theta = (1 - t^2)/1+t^2)\] then the problem is \[[1 - (1 - t^2)/(1 + t^2)]/[1 +(1 - t^2)/(1 + t^2)]\] let \[1 = (1+ t^2)/(1+t^2)\] this will allow for common denominators in the numerator and denominator then \[[( 1+t^2 - 1+t^2)/(1+t^2)]/[( 1 +t^2 + 1 - t^2)/(1 + t^2)]\] this simplifies to \[[2t^2/(1+t^2)]/[2/(1+t^2)]\] dividing by a fraction, flip the denominator and multiply \[2t^2/(1+t^2) \times (1 + t^2)/2 = 2t^2/2\] the answer is t^2 and \[t = \tan(\theta/2)\] so \[t^2 = (\tan(\theta/2))^2\] = \[\tan^2(\theta/2)\]

OpenStudy (anonymous):

Hmm. This doesn't look to simple... Is there an easier way?

OpenStudy (campbell_st):

nope thats it...I wrote all the steps... if you use t = tan(x/2) then cosx = (1 - t^2)/(1+t^2) and sin x = 2t/(1+t^2) its the most painful way to solve trig equations... sometimes called the method of last resort

OpenStudy (anonymous):

Thanks anyway :)

OpenStudy (campbell_st):

if you know sin & cos... its just algebraic manipulation

OpenStudy (anonymous):

hey @order check this method if this is an easy one or not 1-cos x=2sin^2(x/2) 1+cos x=2cos^2(x/2) Substitute them and cancel the two to get tan^2(x/2). if you need these derivations i am ready to help

OpenStudy (anonymous):

I can't seem to get this....

OpenStudy (anonymous):

you know that cosx=cos^2 (x/2)-sin^2 (x/2) therefore cosx=cos^2 (x/2)-1+cos^2 (x/2) (cos^2 (x/2)=1- sin^2 (x/2)) therefore !+cos x=2cos^2 (x/2) got it similarly derive 1-cosx

OpenStudy (anonymous):

then it shud be easy for you in 1-cosx convert the cos terms to sin terms

OpenStudy (anonymous):

I'm sorry... My brain isn't functioning well today.. I'm completely lost! :(

OpenStudy (anonymous):

where are you facing problem?

OpenStudy (anonymous):

From everywhere! From the beginning....

OpenStudy (anonymous):

hpoe you are aware of the formula \[\cos 2x=\cos ^{2}x-\sin^{2} x\]

OpenStudy (anonymous):

from there you derive (1-cosx) how? cos2x=1-sin^2 x-sin^2 x cos2x=1-2sin^x 1-cos2x=2sin^2 x

OpenStudy (anonymous):

if 1-cos 2x=2sin^2 x then 1-cosx=2 sin^2(x/2)